Susan, a single taxpayer, owns and operates a bakery (as a sole proprietorship). The business is not a "specified services" business. In 2019, the business pays $100,000 in W–2 wages, has $150,000 of qualified property, and generates $350,000 of qualified business income. Susan has no other items of income or loss and will take the standard deduction.

Assume the QBI amount is net of the self-employment tax deduction.

What is Susan's qualified business income deduction?
Answer #1

Answer:

Standard deduction in 2019 for Single filers = $12,200

Taxable income (before QBI) = 350000 - 12200 = $337,800

QBI deduction is lesser of:

20% of your qualified business income = 20% * 350000 = $70,000

20% of your taxable income = 20% * 337800 = $67,560

However, since Susan's net income is above threshold value, W2 wage limitations as below also applies:

Greater of:

  • 50% of the W-2 wages paid to employees = 100000 * 50% = $50,000
  • The sum of 25% of such W-2 wages plus 2.5% of the unadjusted basis immediately upon acquisition (UBIA) of the qualified property = 25% * 100000 + 2.5% * 150000 = $28,750

Hence:

Susan's qualified business income deduction = $50,000